40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 1302     Entry time: Fri Aug 16 10:40:20 2013
Author: Erica 
Type: DailyProgress 
Category: fiber optic 
Subject: Spectra Data Locked and Unlocked Laser 

 Tara improved the alignment so we got a little over 1V coming from the fiber alone. We took another run of data of the recombined beam:

laser not locked to cavity:

deltaV = 3.64V

Vmin = 880mV

Vmax = 4.52V

 

laser locked to cavity:

deltaV = 0.17V

Vmax=.236V

Vmin=0.56V

This matches up with the original data taken.

 

We also took data for the noise of the spectrum analyzer.

This can be approximated at a straight line.  I took an average of the points so the noise level is at 1.9*10^-6 Hz/rtHz.

 

Then we took data with the laser locked to the cavity.  The power is much lower because we accidentally misaligned the beam.

As seen in the graph below, the fiber seems to be okay; the plots of the laser unlocked to the cavity match up.

deltaV = 0.17V

Vmax=.236V

Vmin=0.56V

 

We took data for the error signal from the servo. The slope of the error signal for ACAV path is 200kHz/V (see elog entry  1920) .  We are using this to convert from voltage noise to frequency noise.  The shape of the spectrum from the recombined beam follows the shape of the error signal.

0815error.png

 

 

Attachment 2: 0815error.fig  95 kB  Uploaded Fri Aug 16 16:38:38 2013
ELOG V3.1.3-